Suppose you are the manager of Mountain Enterprises, a firm that holds a patent that makes it the exclusive manufacturer of bubble memory chips. Based on the estimates provided by a consultant, you know that the relevant demand and cost functions for bubble memory chips are
Q = 25 - 0.5P;

C = 50 + 5Q.

a. What is the firm's inverse demand function?
b. What is the firm's marginal revenue when producing four units of output?
c. What are the levels of output and price when you are maximizing profits?
d. What will be the level of your profits?

Please show work. Thanks!
Answer #1

♡ function (b) Monopoly Q=25-osp Inverse demand o.sp=25-a p=25-Q ois p=so-aa TR= p.Q = (50-2Q)Q TR a SoQ-2Q² MR =D TR da = soMC = DTC ots(1) Me as Monopoly profit - maximizing Condition MR =MC so-uQ = 5 4Q = sors Q = us Q = 11:25 units substituting iTC = sot 56.25 TC = 106.25 profit - TR-TC = 309.375-106.25 = $203.125